OpenStudy (anonymous):

how would i graph y=4/5x + 14/5

OpenStudy (coconutjj):

y = mx+b m is the slope b is the y-intercept

OpenStudy (anonymous):

y = mx + b m is rise over run so 4 up and 5 over to the right also on the y axis ( your b) is 14/5 so your y intercept is 2.8 point (0,2.8)

OpenStudy (anonymous):

@Austin6i6 so would it look like this?

OpenStudy (nincompoop):

you can graph the equation by first identifying its intercepts - the y-intercept and the x-inceterpt the y-intercept of an equation of a line in the form of y = mx + b is the b such as the case (format) of your equation the x-intercept of an equation of a line may be solved by making the equation to zero, that is: 0 = 4/5 x + 14/5 then solve for the x

OpenStudy (nincompoop):

once you have determined the value of the x and the value of the y understand that x intercept's coordinates is in the format of (x, 0) and the y-intercept's coordinates is in the format of (0, y)

OpenStudy (nincompoop):



OpenStudy (anonymous):

@nincompoop okay that makes a lot more sense thank you!

OpenStudy (nincompoop):

you're welcome